# Quantum Sensing Using Cold Atoms and Molecules for Nuclear and Particle Physics

### v2

@@ V.2 Experimental methods for precision spectroscopy of HCI
Early HCI spectroscopy was performed by classifying the astronomical observations and detecting radiation in fusion researches, including tokamak experiments [243]. These spectroscopic studies primarily focused on the X-ray region and involved mixtures of multiple ionic species. To enable spectroscopy of specific ions, the development of ion sources was essential. The earliest sources of HCIs were based on electron-cyclotron resonance ion sources [244], followed by the invention of electron beam ion sources[245]. Early optical spectroscopy of HCIs was performed in 1987 for fine structure transitions [246], and in 1994 for a hyperfine structure transition [247].
@@
The generated HCIs is extracted from an EBIT and then transferred to a Paul trap. At this stage, the HCIs are hot and must be cooled toward its motional ground state for precision spectroscopy. Due to the absence of broad-linewidth transitions suitable for laser cooling in HCIs, cooling is typically achieved through sympathetic cooling. This method, along with high-precision spectroscopy, has been demonstrated in Ar13+ [249]. In this experiment, an Ar13+ ion is stopped by a Coulomb crystal of Be+, and subsequently laser-cooled using a broad transition in Be+. The experiment is conducted in a cryogenic trap to suppress the BBR shift [250]. The transition at 441 nm serves as the clock transition, with its transition frequency measured to a fractional uncertainty of , referenced to the state-of-the-art 171Yb+ clock. The dominant source of uncertainty in this measurement is the statistical uncertainty, while the total systematic uncertainty for the Ar13+ system is . The primary contributor to systematic uncertainty arises from micromotions, which can be suppressed in the future work [251].
Other ionic species are also investigated, motivated by theoretical proposals. One of the extensively investigated species is Pr ions. Pr9+ was first experimentally examined in Ref.[252]. Although for the state is smaller than ’s listed in Table 5, energies of the long-lived states have been determined with a fractional uncertainty. Pr10+ also has an early spectroscopy report [238]. Combined with an improved theoretical calculation of the energies of the metastable excited states [211], the energy of the state was determined with a fractional uncertainty. Transition frequency measurements were also performed for a narrow-linewidth transition in Ni+12 [253] and many charge states for Xe [254] and Ca [236] ions. The investigations of Xe and Ca ions are motivated by the search for the fifth force through isotope shift measurements (see Sec. VIII). In particular, Ca14+ was recently utilized to set a new constraint on the fifth force[255].
@@ VI Nuclear clock
An even more rigid object against external electric fields than the inner-shell electronic cloud is the nucleus. However, using a typical nuclear transition at keV for precision spectroscopy with a laser is highly challenging. Exceptionally, a low-lying nuclear isomer state in 229Th is barely accessible by laser excitation from the ground state. The existence of this extremely low-energy nuclear excited state was first proposed by Kroger and Reich [256]. The idea of using this isomer state as a clock was initially proposed in 2003 [18]. At the time, the energy of this transition was known with only one-digit precision and was estimated to be lower than its currently known value. However, in recent years, substantial advancements have been made. The transition frequency is now determined to be 2 020 407 384 335(2) kHz, corresponding to an energy of 148.38 nm and 8.3557 eV, with a fractional uncertainty of . Additional details of the transition are summarized in Fig. 7. Note that the next lowest-energy nuclear excited state known today is in 235mU, which is 76.7 eV above the ground state. With this state in 229Th, a clock can be operated based on a nuclear transition. Because the nucleus is times smaller than the atom, it is nearly unaffected by the ac Stark shift.
@@
Historical development of the energy estimates for the first excited state of 229Th is summarized in Table 6. Initially, the energy of the excited state was measured by low-energy gamma rays emitted during the decay of 233U to 229Th [257, 258, 259, 260]. More recent, precise measurements utilized internal conversion and X-ray excitation of the nucleus to a higher excited state than the one involved in the clock transition [261]. In addition, spectroscopy of photons emitted during the relaxation of the clock state further reduced the uncertainty in the transition energy[262], improving the precision of the transition frequency to 4 digit precision. A breakthrough occurred when laser excitation of the transition and subsequent decay of the excited state was observed in 229Th4+ ions doped in a CaF2 crystal [263]. Soon afterthat a similar observation was reported with Th-doped LiSrAlF6 [264]. These studies measured the wavelength with 7 digit precision, confirming that the observation was not crystal specific but resulted from 229Th interacting with the light field. A few months later, spectroscopy using the seventh harmonic generation of a frequency comb determined the transition energy with 2 kHz uncertainty [265]. The observed lifetime of the excited state is 1770(11) s. While this long lifetime provides excellent coherence for long interrogation time, it also makes quick manipulation challenging. To address this, methods to shorten the lifetime were explored [266, 267]. Further studies have also investigated the fine structure [204] and thermal effects [268]. In the field of optical lattice clocks and ion clocks, well-established methodologies can reduce the uncertainty of resonant frequency measurements from 1 kHz to Hz. However, a key difference between these optical clock systems and the 229Th nuclear clock is the lack of a stable continuous-wave laser operating at the resonant frequency of the clock transition.
@@
The nuclear clock can be implemented in two different ways. To prevent the 229Th nucleus from rapidly decaying via internal conversion, the 229Th atoms must be ionized into 229Th3+ ions. These ions can be trapped using an ion trap similar to those used in HCI clocks. This approach has the advantage of leveraging well-established spectroscopy techniques from ion clocks, making it easier to estimate systematic uncertainties. Recently, Laser spectroscopy of 229Th3+ ions in a Paul trap was demonstrated [269]. However, due to the small number of ions in the system, the QPN is large. An alternative method, unique to nuclear clocks, is doping a crystal with 229Th3+ or 229Th4+ ions. Because nucleus is highly isolated from the surrounding electrons, the effects being in a crystal are minimal compared to those on atoms. This approach allows for a large number of ions, [264], to be incorporated, which helps suppress QPN. However, a recent study showed a strong temperature dependence of the transition frequency [268], attributed to changes in the electric field induced by the thermal expansion of the lattice constant.
@@
1976 | <100 | [256] |
1990 | [257] | |
1990 | <5 | [270] |
1994 | 3.5(1.0) | [258] |
2005 | 5.5(1.0) | [271] |
2007 | 7.6(5) | [272] |
2019 | 8.28(17) | [261] |
2019 | 8.30(92) | [259] |
2020 | 8.10(17) | [260] |
2023 | 8.338(24) | [262] |
2024 | 8.367(24) | [273] |
2024 | 8.355 74(3) | [263] |
2024 | 8.355 733(10) | [264] |
2024 | 2 020 407 384 335(2) kHz | [265] |
For fundamental physics applications, the nuclear clock is expected to be highly sensitive to variation of fundamental constants. Because nuclear structure depends on the coupling constant of the strong force , it should be sensitive to variation of [274]. In addition, theoretically predictions suggest that it also exhibits high sensitivity to the variation of , with [204]. Intuitively, this large arises because the small transition energy results from a cancellation between a large repulsive Coulomb energy and an equally large attractive energy from the strong force. The large also makes the nuclear clock a promising tool for searching for ultralight dark matter (see Sec. VII). Another potential research direction involves developing a system that operates without vacuum chamber [275]. Such a system could be advantageous for compact portable optical clock.
@@ VII.1 Theoretical basis
Searches for ultralight dark matter using atomic clocks are sensitive to mass ranges whose the corresponding energy scale is the same as or smaller than the single cycle of the clock operation, which is s. This corresponds to a dark matter mass of eV/c2 or lower. The lower bound for this mass is determined by the overall duration of clock operation. From a cosmological perspective, dark matter is generally expected to have a mass greater than eV in order to account for all observed phenomena with a single candidate. This is because, at this mass, the de Broglie wavelength of dark matter exceeds the size of dwarf galaxies. However, if dark matter consists of multiple components, mass ranges below eV may still be relevant for investigation.
The first proposal for searching for ultralight dark matter using atomic clocks focused on detecting topological dark matter with quadratic coupling to SM particles [299]. This concept initially relied on GPS satellites [300] and required collaboration between multiple research groups across different geographic locations for experimental realization [301]. Later, a new approach was proposed to detect ultralight dilaton dark matter, putting emphasis on the linear coupling to fundamental constants [298]. This proposal included clock comparisons within a single location. After this proposal, various experimental efforts and theoretical analyses of existing data were conducted to put constraints on the ultralight dark matter, starting with a constraint derived from Dy spectroscopy [302].
@@
with G being the Newtonian constant of gravitation and being Planck mass, is the electromagnetic field tensor, is the intensity tensor for the gluon field, is the coupling constant for quantum chromodynamics (QCD), is the function for the running coupling constant for , are the mass of an electron, an up quark and a down quark, respectively, is the anomalous dimension characterizing , and is the spinor for fermions. is integer, and correspond to linear and quadratic coupling, respectively. Some studies parametrize the coupling constant as , in which case is associated with the energy scale for new physics [299, 304]. Each term in Eq. VII.1 describes a tree-level interaction, where a SM particle emits an ultralight dark matter particle, with the coupling constant for this interaction being . Well-known particles that satisfy these properties are the dilaton and moduli. Similar interactions can be assumed for axions when CP symmetry is violated [305].
When couples to ordinary matter, fundamental constants vary according to the coupling constants in the Lagrangian in the following way.
@@
Searching for ultralight dark matter through this interaction is relatively easy to implement in a laboratory because atomic transitions and the size of an atom depend differently on fundamental constants [306, 307]. Take as an example. The wavelength of a narrow-linewidth laser used for precision spectroscopy is stabilized by locking it to a stable reference cavity, where the mirror separation is determined by the spacer material. Since the laser wavelength is proportional to , the frequency scales as , where is the Bohr radius. However, an atomic transition does not generally follow the same -dependence, as discussed in Sec. IV. This difference in -scaling leads to a periodic variation in the atomic resonance relative to the cavity resonance when couples to the ultralight dark matter field .
@@ VII.2 Status of experimental searches
At the end of an experimental analysis, the amplitudes of a sinusoidal oscillation in the frequency ratio are obtained. The frequency assumed for the sinusoidal fit can be arbitrarily chosen within the range up to the Nyquist frequency. As a result, a single dataset for measured over time can provide constraints on dilaton ultralight dark matter across a wide mass range, as shown in Fig. 8. So far, no experimental results detected sinusoidal oscillations with amplitudes significantly larger than their statistical uncertainties. The resulting constraints exclude coupling constants stronger than a certain threshold for a given mass. In Fig. 8, these excluded regions are represented as shaded area. In this kind of searches for new particles, it is common practice to plot mass on the horizontal axis and a coupling constant for a specific interaction on the vertical axis.
@@
Comparisons between optical clocks have no sensitivity to , primarily because all optical transitions are more or less proportional to , resulting in a constant when frequency ratio is taken. Instead, comparisons between an electronic transition or a reference cavity and a microwave transition between hyperfine levels exhibit a high sensitivity. The comparison between an Si cavity and a hydrogen maser [309] reported constraints down to eV. The lower bound on mass is limited by the measurement duration. A long-term comparison between an Yb optical lattice clock and a Cs fountain clock, spanning 298 days, provided constraints for smaller masses [308].
For and , the contribution of clock comparisons is relatively small compared to satellite [315, 295] and torsion balance [316, 295] experiments, as these macroscopic experiments involve much large masses. Although Fig. 8(c) and (d) show slight excess in the performance of the clock comparisons at low masses, the final result of the MICROSCOPE experiment [314], which improved by a factor of 4.6 compared to the initial result shown in the plot, is not yet included. With this updated result, most clock comparison constraints become marginal. Note that the constraint by Yb does not follow the formalism discussed in the previous subsection but arises from modulation in nuclear charge radii [312].
@@ VIII.1 Theoretical basis
Isotope shifts refer to the frequency difference of an atomic transition between isotopes. This phenomenon was already known in the first half of the 20th century [330, 331]. For optical transitions with THz, isotope shifts typically range from 10 MHz to 1 GHz. These shifts arise mainly from two sources and can be phenomenologically described as:
The first term, referred to as the mass shift, results from differences in nuclear mass. The second term, known as the field shift, is caused by variation in the nuclear charge radius due to differences in the number of neutrons. In Eq. 10 and the following discussion, the subscript labels a transition, and , unless otherwise specified. For example, is the transition frequency for the transition of the isotope with mass number . is the inverse mass of the isotope with mass number , is the th moment of the charge radius for the isotopes with mass number , and and are constants characterizing the mass shift and field shift, respectively. can be eliminated two isotope shift equations for transitions and .
Here, . The equation shows that is a linear function of . The two-dimensional plot of versus is called the King plot [332].
@@
Higher-order effects of the nuclear charge radius, which is within the SM, also induce nonlinearity. The following equation is typically assumed in the state-of-the-art analysis.
The third and fourth terms correspond to the 4th-moment shift and the quadratic field shift, respectively, where is the second order perturbative effect due to . The equation analogous to Eq. 11 is
In this equation, the third, fourth, and fifth terms all contribute to nonlinearity. As a result, properly subtracting background effects from the SM contributions are crucial when searching for new forces.
@@
Several conditions are necessary for high-sensitivity measurements. Searching for new forces through King plot nonlinearity requires an element with at least four stable isotopes. The more isotopes and transitions available, the more information can be extracted. Narrow-linewidth transitions are preferred for high-precision measurements. In addition, spinless nuclei are favored to avoid spin-spin interactions that are difficult to precisely estimate. In fact, significant nonlinearity is observed when isotopes with nonzero spin are included in the King plot for Sr [338] and Yb [122], with the second-order hyperfine interaction fully explaining the observed nonlinearity in Sr [339]. For higher sensitivity, transitions involving excited states with large difference in radial wave functions are preferable. For example, a combination of and transitions are expected to yield higher sensitivity in the King plot compared to a combination of and the transitions. To suppress unwanted backgrounds arising from SM effects, light nuclei are preferred, because of their small charge radii and reduced relativistic effects. One key source of uncertainty is , which can be eliminated by using three narrow-linewidth transitions [340]. Another approach to analyzing isotope shift data is to sort them by isotopes rather than transitions[341], which claimed to provide constraints on the new bosons comparable to those obtained through conventional King plot analysis.
This new force coupling to an electron and a neutron is introduced phenomenologically, without necessarily being supported by strong theoretical motivations. However, one benefit of this approach is its sensitivity to the X17 particle [342]. The force discussed in this section is spin-independent, while spin-dependent fifth force searches are covered in a recent review paper [13]. Also, the formalism and recent progresses in the King plot analysis is also summarized in another recent review paper[343].

### v3

@@ V.2 Experimental methods for precision spectroscopy of HCI
Early HCI spectroscopy was performed by classifying the astronomical observations and detecting radiation in fusion researches, including tokamak experiments [258]. These spectroscopic studies primarily focused on the X-ray region and involved mixtures of multiple ionic species. To enable spectroscopy of specific ions, the development of ion sources was essential. The earliest sources of HCIs were based on electron-cyclotron resonance ion sources [259], followed by the invention of electron beam ion sources[260]. Early optical spectroscopy of HCIs was performed in 1987 for transitions between fine structures [261], and in 1994 for a transition between a hyperfine structure [262].
@@
The generated HCIs are extracted from an EBIT and then transferred to a Paul trap. At this stage, the HCIs are hot and must be cooled toward its motional ground state for precision spectroscopy. Due to the absence of broad-linewidth transitions suitable for laser cooling in HCIs, cooling is typically achieved through sympathetic cooling. This method, along with high-precision spectroscopy, has been demonstrated in Ar13+ [264]. In this experiment, an Ar13+ ion is stopped by a Coulomb crystal of Be+, and subsequently laser-cooled using a broad transition in Be+. The experiment is conducted in a cryogenic trap to suppress the BBR shift [265]. The transition at 441 nm serves as the clock transition, with its transition frequency measured to a fractional uncertainty of , referenced to the state-of-the-art 171Yb+ clock. The dominant source of uncertainty in this measurement is the statistical uncertainty, while the total systematic uncertainty for the Ar13+ system is . The primary contributor to systematic uncertainty arises from micromotions, which can be suppressed in the future work [266].
Other ionic species are also investigated, motivated by theoretical proposals. One of the extensively investigated species is Pr ions. Pr9+ was first experimentally examined in Ref.[267]. Although for the state is smaller than ’s listed in Table 5, energies of the long-lived states have been determined with a fractional uncertainty. Pr10+ also has an early spectroscopy report [253]. Combined with an improved theoretical calculation of the energies of the metastable excited states [223], the energy of the state was determined with a fractional uncertainty. Transition frequency measurements were also performed for a narrow-linewidth transitions in Ni12+ [268, 269] and many charge states for Xe [270], Ca [251], and Os [271] ions. An observed narrow-linewidth transition in Ni12+ has an expected linewidth of 8 mHz, and the search was performed by a close collaboration between a theoretical prediction of the transition energy and an experimental search for it [269]. The investigations of Xe, Ca, and Os ions are motivated by the search for a fifth force through isotope shift measurements (see Sec. VIII). In particular, Ca14+ was recently utilized to set a new constraint on the fifth force[272].
@@ VI Nuclear clock
An object even more rigid against external electric fields than the inner-shell electronic cloud is the nucleus. However, using a typical nuclear transition at keV for precision spectroscopy with a laser is extremely challenging. Exceptionally, a low-lying nuclear isomer state in 229Th is barely accessible by laser excitation from the ground state. The existence of this extremely low-energy nuclear excited state was first proposed by Kroger and Reich [273]. The idea of using this isomer state as a clock was initially proposed in 2003 [18]. At the time, the energy of this transition was known with only one-digit precision and was estimated to be lower than its currently known value. However, in recent years, substantial advancements have been made. The transition frequency is now determined to be 2 020 407 384 335(2) kHz, corresponding to a wavelength of 148.38 nm and energy of 8.3557 eV, with a fractional uncertainty of . Additional details of the transition are summarized in Fig. 7. Note that the next lowest-energy nuclear excited state known today is in 235mU, which is 76.7 eV above the ground state. With this state in 229Th, a clock can be operated based on a nuclear transition. Because the nucleus is times smaller than the atom, it is nearly unaffected by the ac Stark shift.
@@
Historical development of the energy estimates for the first excited state of 229Th is summarized in Table 6. Initially, the energy of the excited state was measured by low-energy gamma rays emitted during the decay of 233U to 229Th [274, 275, 276, 277]. More recent, precise measurements utilized internal conversion and x-ray excitation of the nucleus to a higher excited state than the one involved in the clock transition [278]. In addition, spectroscopy of photons emitted during the relaxation of the clock state further reduced the uncertainty in the transition energy[279], improving the precision of the transition frequency to four digits. A breakthrough occurred when laser excitation of the transition and subsequent decay of the excited state was observed in 229Th4+ ions doped in a CaF2 crystal [280]. Soon after that a similar observation was reported with Th-doped LiSrAlF6 [281]. These studies measured the wavelength with seven digit precision, confirming that the observation was not crystal specific but resulted from 229Th interacting with the light field. A few months later, spectroscopy using the seventh harmonic generation of a frequency comb determined the transition energy with 2 kHz uncertainty [282]. The observed lifetime of the excited state is 1770(11) s. In the field of optical lattice clocks and ion clocks, well-established methodologies can reduce the uncertainty of resonant frequency measurements from 1 kHz to Hz. However, a key difference between these optical atomic clock systems and the 229Th nuclear clock is the lack of a stable continuous-wave laser operating at the resonant frequency of the clock transition.
With an observation of the transition and the determination of the transition frequency, the investigation is moved to the further characterization of the transition. While this long lifetime provides excellent coherence for long interrogation time, it also makes quick manipulation challenging. To address this, methods to shorten the lifetime were explored, using laser-induced quenching [283] and a thin film of the host material [284]. Further reduction of the lifetime of the isomer state is achieved by performing spectroscopy in a material opaque to the light resonant to the nuclear clock transition [285]. ThO2 has a band gap of 6 eV, smaller than the nuclear transition at 8.4 eV. This allows decays through internal conversion, reducing the lifetime to s. Further studies have also investigated thermal effects [286] and reproducibility in the transition frequency between different batches of doped crystals at the temperature where the first order thermal sensitivity is zero [287]. A theoretical analysis pointed out that the transition frequency for the nuclear clock transition can be different between host materials, ions with different charge, and bare nuclei [288].
@@
The nuclear clock can be implemented in two different ways. To prevent the 229Th nucleus from rapidly decaying via internal conversion, the 229Th atoms must be ionized into 229Th3+ ions. These ions can be trapped using an ion trap similar to those used in HCI clocks. This approach has an advantage of leveraging well-established spectroscopy techniques from ion clocks, making it easier to estimate systematic uncertainties. However, due to the small number of ions in the system, the QPN is large. Towards this direction, laser spectroscopy of 229Th3+ ions in a Paul trap was demonstrated [289]. Another reports sympathetically cools Th3+ ions with 88Sr+ ions [290], and performed spectroscopy on hyperfine structures and isotope shifts in Th3+ ions [291]. An alternative method, unique to nuclear clocks, is doping a crystal with 229Th3+ or 229Th4+ ions. Because nucleus is highly isolated from the surrounding electrons, the effects being in a crystal are minimal compared to those on atoms. This approach allows for a large number of ions, e.g., [281], to be incorporated, which helps suppress QPN. However, a recent study showed a strong temperature dependence of the transition frequency [286], attributed to changes in the electric field induced by the thermal expansion of the lattice constant, unless temperature is set at a specific temperature where the first order thermal shift is zero [287].
@@
| 1976 | <100 | [273] |
| 1990 | [274] | |
| 1990 | <5 | [292] |
| 1994 | 3.5(1.0) | [275] |
| 2005 | 5.5(1.0) | [293] |
| 2007 | 7.6(5) | [294] |
| 2019 | 8.28(17) | [278] |
| 2019 | 8.30(92) | [276] |
| 2020 | 8.10(17) | [277] |
| 2023 | 8.338(24) | [279] |
| 2024 | 8.367(24) | [295] |
| 2024 | 8.355 74(3) | [280] |
| 2024 | 8.355 733(10) | [281] |
| 2024 | 2 020 407 384 335(2) kHz | [282] |
For fundamental physics applications, the nuclear clock is expected to be highly sensitive to variation of fundamental constants. Because nuclear structure depends on the coupling constant of the strong force , it should be sensitive to variation of [296]. In addition, theoretical predictions suggest that it also exhibits high sensitivity to the variation of , with [216]. Intuitively, this large arises because the small transition energy results from a cancellation between a large repulsive Coulomb energy and an equally large attractive energy from the strong force. The large also makes the nuclear clock a promising tool for searching for ultralight dark matter (see Sec. VII). Another potential research direction involves developing a system that operates without vacuum chamber [297]. Such a system could be advantageous for compact portable optical atomic clock.
@@ VII.1 Theoretical basis
Searches for ultralight dark matter using atomic clocks are sensitive to mass ranges whose energy scale is the same as or smaller than the single cycle of the clock operation, which is s. This corresponds to a dark matter mass of eV/c2 or lower. The lower bound for this mass is determined by the overall duration of clock operation. From a cosmological perspective, dark matter is generally expected to have a mass greater than eV in order to account for all observed phenomena with a single candidate. This is because the de Broglie wavelength of dark matter exceeds the size of dwarf galaxies at this mass. However, if dark matter consists of multiple components, mass ranges below eV may still be relevant for investigation.
The first proposal for a search for ultralight dark matter using atomic clocks focused on detecting topological dark matter with quadratic coupling to SM particles [321]. This concept initially relied on GPS satellites [322] and required collaboration between multiple research groups across different geographic locations for experimental realization [323]. Later, a new approach was proposed to detect ultralight dilaton dark matter, putting emphasis on the linear coupling to fundamental constants [320]. This proposal included clock comparisons within a single location. After this proposal, various experimental efforts and theoretical analyses of existing data were conducted to put constraints on the ultralight dark matter, starting with a constraint derived from Dy spectroscopy [324].
@@
Here, with G being the Newtonian constant of gravitation and being Planck mass, is the electromagnetic field tensor, is the intensity tensor for the gluon field, is the coupling constant for quantum chromodynamics (QCD), is the function for the running coupling constant for , are the mass of an electron, an up quark and a down quark, respectively, is the anomalous dimension characterizing , and is the spinor for fermions. is integer, and correspond to linear and quadratic coupling. Some studies parametrize the coupling constant as , in which case is associated with the energy scale for new physics [321, 326]. Each term in Eq. VII.1 describes a tree-level interaction, where a SM particle emits an ultralight dark matter particle, with the coupling constant for this interaction being . Well-known particles that satisfy these properties are the dilaton and moduli. Similar interactions can be assumed for axions when CP symmetry is violated [327].
When couples to ordinary matter, fundamental constants vary according to the coupling constants in the Lagrangian in the following way:
@@
Searching for ultralight dark matter through this interaction is relatively easy to implement in a laboratory because transition frequencies of atomic transitions and the size of an atom depend differently on fundamental constants [328, 329]. Take as an example. The wavelength of a narrow-linewidth laser used for precision spectroscopy is stabilized by locking it to a stable reference cavity, where the mirror separation is determined by the spacer material. Since the laser wavelength is proportional to , the frequency scales as , where is the Bohr radius. However, an atomic transition does not generally follow the same -dependence, as discussed in Sec. IV. This difference in -scaling leads to a periodic variation in the atomic resonance relative to the cavity resonance when couples to the ultralight dark matter field .
@@ VII.2 Status of experimental searches
At the end of an experimental analysis, the amplitudes of a sinusoidal oscillation in the frequency ratio are obtained. The frequency assumed for the sinusoidal fit can be arbitrarily chosen within the range up to the Nyquist frequency. As a result, a single dataset for measured over time can provide constraints on dilaton ultralight dark matter across a wide mass range, as shown in Fig. 8. So far, no experimental results detected sinusoidal oscillations with amplitudes significantly larger than their statistical uncertainties. The resulting constraints exclude coupling constants stronger than a certain threshold for a given mass. In Fig. 8, these excluded regions are represented as shaded areas. In this kind of searches for new particles, it is common practice to plot mass on the horizontal axis and a coupling constant for a specific interaction on the vertical axis.
@@
Comparisons between optical atomic clocks have no sensitivity to , primarily because all transition frequencies for optical transitions are more or less proportional to , resulting in a constant when the frequency ratio is taken. Instead, comparisons between an electronic transition or a reference cavity and a microwave transition between hyperfine levels exhibit a high sensitivity. The comparison between an Si cavity and a hydrogen maser [331] reported constraints down to eV. The lower bound on mass is limited by the measurement duration. A long-term comparison between an Yb optical lattice clock and a Cs fountain clock, spanning 298 days, provided constraints for smaller masses [330].
For and , the contribution of clock comparisons is relatively small compared to satellite [337, 317] and torsion balance [338, 317] experiments, as these macroscopic experiments involve much larger masses. Although Fig. 8(c) and 8(d) show a slight excess in the performance of the clock comparisons at low masses, the final result of the MICROSCOPE experiment [336], which improved by a factor of 4.6 compared to the initial result shown in the plot, is not yet included. With this updated result, most clock comparison constraints become marginal. Note that the constraint by Yb does not follow the formalism discussed in the previous subsection but arises from modulation in nuclear charge radii [334].
@@ VIII.1 Theoretical basis
Isotope shifts refer to the frequency difference of an atomic transition between isotopes. This phenomenon was already known in the first half of the 20th century [352, 353]. For optical transitions with THz, isotope shifts typically range from 10 MHz to 1 GHz. These shifts arise mainly from two sources and can be phenomenologically described as
The first term, referred to as the mass shift, results from differences in nuclear mass. The second term, known as the field shift, is caused by variation in the nuclear charge radius due to differences in the number of neutrons. In Eq. 9 and the following discussion, the subscript labels a transition, and , unless otherwise specified. For example, is the transition frequency for the transition of the isotope with mass number . is the inverse mass of the isotope with mass number , is the th moment of the charge radius for the isotopes with mass number , and and are constants characterizing the mass shift and field shift, respectively. can be eliminated with two isotope shift equations for transitions and .
Here, . The equation shows that is a linear function of . The two-dimensional plot of vs is called the King plot [354].
@@
Higher-order effects of the nuclear charge radius, which is within the SM, also induce nonlinearity. The following equation is typically assumed in the state-of-the-art analysis:
The third and fourth terms correspond to the fourth-moment shift and the quadratic field shift, where is the second-order perturbative effect due to . The equation analogous to Eq. 10 is
In this equation, the third, fourth, and fifth terms all contribute to nonlinearity. As a result, properly subtracting background effects from the SM contributions are crucial when searching for the new forces.
@@
Several conditions are necessary for high-sensitivity measurements. Searching for new forces through King plot nonlinearity requires an element with at least four stable isotopes. The more isotopes and transitions available, the more information can be extracted. Narrow-linewidth transitions are preferred for high-precision measurements. In addition, spinless nuclei are favored to avoid spin-spin interactions that are difficult to estimate precisely. In fact, significant nonlinearity is observed when isotopes with nonzero spin are included in the King plot for Sr [360] and Yb [125], with the second-order hyperfine interaction fully explaining the observed nonlinearity in Sr [361]. For higher sensitivity, transitions involving excited states with large difference in radial wave functions are preferable. For example, a combination of and transitions is expected to yield higher sensitivity in the King plot compared to a combination of and the transitions. To suppress unwanted backgrounds arising from SM effects, light nuclei are preferred, because of their small charge radii and reduced relativistic effects. One key source of uncertainty is , which can be eliminated by using three narrow-linewidth transitions [362]. Another approach to analyzing isotope shift data is to sort them by isotopes rather than transitions[363], which claimed to provide constraints on the new bosons comparable to those obtained through conventional King plot analysis.
This new force coupling to an electron and a neutron is introduced phenomenologically, without necessarily being supported by strong theoretical motivations. However, one benefit of this approach is its sensitivity to the X17 particle [364]. The force discussed in this section is spin-independent, while spin-dependent fifth force searches are covered in a recent review paper [13]. The formalism and recent progress in the King plot analysis are also summarized in another recent review paper[365].